OpenStudy (anonymous):

Which is a counterexample that disproves the conjecture? A student concludes that if x is a real number, then x ≤ x3. A. 1 B. 0 C. -1/2 D. -2

OpenStudy (anonymous):

A counter example is giving an example that disproves what was states so we need to to disprove that for any x where x is a real number then \(x \leq x^3 \) So lets find a value that is a real number buttttttt \(x>x^3\) So lets just plug in a couple of points to figure this out so lets x=1 then x^3=1 so 1=1 so that wont disprove it Let x=2 then x^3=8 and 2<8 so this also does not disprove How abt we take a negative number Let x=-2 then x^3=-8 and -2>-8 SOOOO VOILAAA we have found a value of x that x>x^3 So this disproves the conjecture